OpenStudy (anonymous):

The average of 5 consecutive even numbers A,B,C,D and E is 52. What is the product of B and E?

OpenStudy (anonymous):

i think it could be like this x = first number A x+2 = B x+4 = C x+6 = D x+8= E x+x+2+x+4+x+6+x+8 = 52

OpenStudy (anonymous):

What is the product of B & E

OpenStudy (anonymous):

sorry (x+x+2+x+4+x+6+x+8 )/5 = 52

OpenStudy (anonymous):

u need to solve for x first

OpenStudy (anonymous):

ok really very much thanks i am very proud of you